OpenStudy (anonymous):

What is 120% of 185?

OpenStudy (anonymous):

Once again, plug in: %= x/100 of = times is ~ =

OpenStudy (anonymous):

37

OpenStudy (anonymous):

Eg What is 120% of 185? x = 120/100 times 185?